OpenStudy (anonymous):

please help me factor this: 3(x+2)^2-10(x+2)+7

hero (hero):

Hint: Expand it first

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

no

hero (hero):

Well, what did you get?

OpenStudy (anonymous):

idk the steps from there

OpenStudy (anonymous):

subtitution, t=x+2

hero (hero):

Yeah, that too

OpenStudy (anonymous):

i don't even know where to start with this problem

hero (hero):

Either way works. t = x+ 2 is a shortcut

OpenStudy (anonymous):

3t^2-10t+7=(3t-1)(3t-7)

hero (hero):

The only difference is, when you do it that way, you have to substitute x + 2 back in

hero (hero):

Which @love_math hasn't done yet

OpenStudy (anonymous):

ohh okay thanks guys i appreciate your help

hero (hero):

(3(x+2)-1)(3(x+2)-7)

hero (hero):

(3x+6-1)(3x+6-7) = (3x+5)(3x-1)

hero (hero):

That is the correct one